# ]Ling Team, Ant Group

### v1

@@ 1 Introduction
Building upon these observations, we derive a unified scaling law for efficiency leverage of MoE. This formula integrates the combined effects of the compute budget, activation ratio, and expert granularity. It enables us to directly predict an MoE architecture’s EL for a given configuration, providing principled guidance for efficient MoE architectural design. As a practical demonstration, Figure 1 illustrates the estimated EL under compute budgets of $201e$ 20 and $221e$ 22 FLOPs.
According to our derived scaling law for EL, we predict that an MoE model with a 3.1% activation ratio and an expert granularity of 12 will achieve over 7x computational efficiency under a $221e$ 22 FLOPs compute budget. To empirically validate this, we designed and trained “Ling-mini-beta,” a pilot model for the Ling-2.0 series, with 0.85B activated parameters and 17.5B total parameters. The model was trained on a 1T-token high-quality dataset and benchmarked against its dense counterpart, a 6.1B parameter model. Experimental results show that when trained on the same 1T-token dataset, Ling-mini-beta achieves a lower final training loss and exhibits a slight performance advantage across a suite of downstream tasks. This outcome confirms our theoretical prediction, validating that this MoE architecture yields an efficiency gain of over 7x. These findings provide a solid theoretical and empirical foundation for the design of future large-scale, efficient MoE models.
(a) FLOPs = $201e$ 20
(b) FLOPs = $221e$ 22
Figure 1: Estimated MoE efficiency leverage using Eq. (13) for FLOPs budgets of $201e$ 20 and $221e$ 22.
@@ 2.2 Scaling Laws for MoE Optimal Hyper-parameters
The performance of a MoE model is sensitive to its hyperparameters. To ensure that our subsequent architectural comparisons are reliable, it is crucial to evaluate each configuration under its optimal hyperparameter settings. Therefore, we first conduct a preliminary study to establish the scaling laws for optimal MoE hyperparameters. Previous research (Bi et al., 2024) has established that the optimal hyperparameters are primarily a function of the total computational budget. Accordingly, we performed a hyperparameter search across a compute range of $181e$ 18 to $203e$ 20 FLOPs, using a Warmup-Stable-Decay (WSD) learning rate schedule (Hu et al., 2024). We trained multiple models, varying both learning rate and batch size, which were sampled from a log-base-2 grid. Specifically, the exponents for the learning rate ranged from -11 to -9.0, and for the batch size, from 18 to 21. To make this analysis tractable, we initially fixed the MoE configuration to one with 64 experts, of which 4 are activated per token, plus an additional shared expert (resulting in an activation ratio $A=7.8\%$ and a granularity $G=2$). Detailed settings of the experimental models are available in the Appendix B. We then verified that the conclusions from this configuration generalize across different activation ratios.
@@
(a) $A$ = 10.9%
(b) $A$ = 7.8%
(c) $A$ = 4.7%
Figure 3: Validation of MoE hyperparameters scaling laws across different activation ratios ($A$). “Near-optimal” refers to hyperparameters achieving a loss within 0.25% of the optimal ones.
To validate the generalizability of these laws, we conduct experiments on MoE models with varying activation ratios. We used the derived laws to predict optimal hyperparameters at a compute budget of $203e$ 20 FLOPs, after fitting them on data up to $201e$ 20 FLOPs. As shown in Figure 3, the predicted optimal regions effectively capture the best-performing hyperparameters for activation ratios from 4.7% to 10.9%, demonstrating that the laws can be applied to MoE models within this range of activation rates. This confirms that our hyperparameter scaling laws provide a reliable foundation for exploring diverse MoE architectures under fair and near-optimal training conditions.
@@ 2.3 Scaling laws for MoE Optimal Model-Data Allocation
MoE $M^{\text{opt}}=0.3187\cdot C^{0.4975}$ $D^{\text{opt}}=3.1382\cdot C^{0.5025}$
@@ 3 Efficiency Leverage: Metric for Quantifying MoE Compute-Efficiency
[…]
@@
To make the analysis tractable, we assume the effects of these dimensions are largely independent and conduct systematic ablation studies. We start with a baseline MoE architecture (2 of 64 experts activated, plus one shared expert) and vary one dimension at a time across a range of compute budgets (from $183e$ 18 to $203e$ 20 FLOPs). To ensure a fair and robust comparison, we leverage the findings from our preliminary studies (Sections 2.2 and 2.3). For each architecture and compute budget, we determine the reasonable model size ($M$) and data size ($D$) using our derived allocation laws and configure training with optimal hyperparameters from our hyperparameter scaling laws. This rigorous protocol ensures that each architecture is evaluated at or near its peak potential for a given budget, yielding reliable and cost-effective conclusions. Further details on the experimental setup are provided in Appendix B. Next, we first empirically analyze the impact of each dimension on EL, and then integrate our empirical findings into a unified scaling law to models the relationship between MoE configurations and the resulting EL.
@@ 4.1.1 Optimal Expert Activation Ratio
The IsoFLOPs curves, presented in Figure 5(a), reveal a clear and consistent trend. Across all tested FLOPs budgets (from $181e$ 18 to $203e$ 20), loss monotonically decreases with activation ratio, following a power-law pattern. For all configurations, the lowest tested ratio of 0.8% consistently yields the minimum loss. This finding suggests a core principle: for a fixed computational cost, greater model sparsity (i.e., lower activation ratio) leads to higher parameter efficiency.
@@
(a) IsoFLOPs curves for varying $A$
(b) Loss and EL scaling curve over varying $A$.
Figure 5: Impact of the Activation Ratio $A$ on Loss and Efficiency. (a) At any fixed compute budget (each colored line), lower activation ratios yield lower loss. The orange stars mark the optimal (lowest) loss point. (b) Loss and EL scaling curves illustrate that EL increases with both higher compute budgets and lower activation ratios, showing that MoE advantages are magnified at scale.
@@ 4.1.3 Optimal Shared Expert Ratio
where $S^{\text{opt}}$ is the optimal shared expert that minimizes the training loss $\mathcal{L}$ under a fixed FLOPs budget $C$, model size $M$, activation ratio $A$, and granularity $G$. Our experiments, as depicted in Figure 7(a), reveal a U-shaped relationship between the shared expert ratio and training loss. The minimum loss is generally achieved at a relatively low shared expert ratio, while having no shared experts ($S=0\%$) usually results in suboptimal performance. Furthermore, we observe a subtle trend where the optimal sharing ratio appears to scale with the compute budget. This is supported by our empirical scaling law (EL) analysis in Figure 7(b), which shows that lower FLOPs budgets ($\leq 10^{20}$) benefit from a slightly higher sharing ratio ($S=16.7\%$), whereas larger budgets ($>10^{20}$) achieve greater efficiency with a lower ratio ($S=8.3\%$).
@@
Figure 7: Impact of the Shared Ratio $S$ on Loss and Efficiency. (a) Loss curves demonstrate that a low, non-zero sharing ratio minimizes training loss, outperforming both no shared experts ($S=0\%$) and highly shared configurations.. (b) EL analysis reveal that the optimal sharing ratio is higher ($S=16.7\%$) for smaller FLOPs ($<10^{20}$) and decreases to $S=8.3\%$ for larger FLOPs ($>10^{20}$).
@@ 4.1.4 Other Configurations of MoE Architecture
First, we analyze replacing the initial MoE layers with dense layers while keeping total FLOPs constant (e.g., 60-layer models with the first 1–3 layers set dense). The experimental results show that replacing the first few layers with dense layers has a minor impact on performance, with efficiency leverage close to 1 within a FLOPs budget of up to $3e{20}$ FLOPs. This adjustment reduces the total number of parameters and mitigates routing imbalances, making it a valuable design optimization. As FLOPs budgets increase, the optimal dense proportion also grows; for example, at $181e$ 18 FLOPs, the optimal dense proportion is zero. As the compute budget increases to $203e$ 20 FLOPs, the optimal dense layer proportion shifts to approximately $2/60$ or $3/60$.
@@ 4.2.1 Separable Scaling Laws for Efficiency Leverage
(a) Scaling with Activation Ratio ($A$)
@@
Figure 8: Scaling Behavior of Efficiency Leverage (EL). (a) With fixed granularity ($G=2$), EL follows a power law with respect to $A$ across all tested compute budgets ($C$). (b) With a fixed activation ratio ($A=3.1\%$), EL’s scaling with $G$ conforms to a log-polynomial law across all compute budgets. (c) With both activation ratio ($A$) and granularity ($G$) held constant, EL scales with compute according to a standard power law.
@@ 5 Ling-mini-beta: More Efficient MoE Language Model
Based on the findings in Section 4, we identify the efficient architectural configuration within the current MoE framework. To validate the effectiveness of this configuration, we design an MoE model with 0.855 billion active parameters out of a total of 17.5 billion (referred to as “Ling-mini-beta,” a pilot model for Ling-2.0 series) and test it using 1 trillion tokens of training data. Ling-mini-beta is configured with a granularity $G$ of 12 and an activation ratio $A$ of just 3.4%. Referring to Figure 1, at the $221e$ 22 FLOPs compute budget, we hypothesize that Ling-mini-beta achieves more than 7× in compute-efficiency leverage over a comparable dense model. Concurrently, we train a traditional dense model with 6.1 billion parameters (named “Dense-6.1B”) for comparison. This section presents a detailed analysis of the performance differences between Ling-mini-beta and the conventional dense model Dense-6.1B, highlighting that the active parameter count, training costs, and downstream inference costs of Dense-6.1B are more than seven times those of Ling-mini-beta.
@@ 5.1 Model and Training Details
  • Training Setting. Both Ling-mini-beta and Dense-6.1B were trained using the AdamW optimizer (Loshchilov and Hutter, 2017) with hyperparameters set as follows: $\beta_{1}=0.9$, $\beta_{2}=0.95$, and weight decay of 0.1. Gradient clipping norm is set to 1.0. The learning rate schedule employs a WSD (warmup-stable-decay) strategy (Hu et al., 2024). According to the hyperparameter scaling laws for dense and MoE models, the maximum learning rates were set to $3.78e{-4}$ for Ling-mini-beta and $2.93e{-4}$ for Dense-6.1B. The batch sizes were configured as 1792 and 2048, respectively.
@@
Model $n_{layers}$ $d_{model}$ $d_{ffn}$ $d_{expert}$ $n_{heads}$ $n_{kv\_head}$ $E$ $E_{a}$ $E_{s}$ $N$ $N_{a}$
@@ Appendix A Notation
$N$ Number of non-vocabulary parameters.
@@ Appendix F List of Experimental Models
$n{layers}$ $d_{model}$ $n_{\text{heads}}$ $E$ $E_{a}$ $E_{s}$ $d_{\text{expert}}$ $B$ $\eta$ Max training FLOPs

### v2

@@ 1 Introduction
Building upon these observations, we derive a unified scaling law for efficiency leverage of MoE. This formula integrates the combined effects of the compute budget, activation ratio, and expert granularity. It enables us to directly predict an MoE architecture’s EL for a given configuration, providing principled guidance for efficient MoE architectural design. As a practical demonstration, Figure 1 illustrates the estimated EL under compute budgets of $1e20$ and $1e22$ FLOPs.
According to our derived scaling law for EL, we predict that an MoE model with a 3.1% activation ratio and an expert granularity of 12 will achieve over 7x computational efficiency under a $1e22$ FLOPs compute budget. To empirically validate this, we designed and trained “Ling-mini-beta,” a pilot model for the Ling-2.0 series, with 0.85B activated parameters and 17.5B total parameters. The model was trained on a 1T-token high-quality dataset and benchmarked against its dense counterpart, a 6.1B parameter model. Experimental results show that when trained on the same 1T-token dataset, Ling-mini-beta achieves a lower final training loss and exhibits a slight performance advantage across a suite of downstream tasks. This outcome confirms our theoretical prediction, validating that this MoE architecture yields an efficiency gain of over 7x. These findings provide a solid theoretical and empirical foundation for the design of future large-scale, efficient MoE models.
(a) FLOPs = $1e20$
(b) FLOPs = $1e22$
Figure 1: Estimated MoE efficiency leverage using Eq. (13) for FLOPs budgets of $1e20$ and $1e22$.
@@ 2.2 Scaling Laws for MoE Optimal Hyper-parameters
The performance of a MoE model is sensitive to its hyperparameters. To ensure that our subsequent architectural comparisons are reliable, it is crucial to evaluate each configuration under its optimal hyperparameter settings. Therefore, we first conduct a preliminary study to establish the scaling laws for optimal MoE hyperparameters. Previous research (Bi et al., 2024) has established that the optimal hyperparameters are primarily a function of the total computational budget. Accordingly, we performed a hyperparameter search across a compute range of $3e17$ to $3e20$ FLOPs, using a Warmup-Stable-Decay (WSD) learning rate schedule (Hu et al., 2024). We trained multiple models, varying both learning rate and batch size, which were sampled from a log-base-2 grid. Specifically, the exponents for the learning rate ranged from -11 to -9.0, and for the batch size, from 18 to 21. To make this analysis tractable, we initially fixed the MoE configuration to one with 64 experts, of which 4 are activated per token, plus an additional shared expert (resulting in an activation ratio $A=7.8\%$ and a granularity $G=2$). Detailed settings of the experimental models are available in the Appendix B. We then verified that the conclusions from this configuration generalize across different activation ratios.
@@
(a) AA = 10.9%
(b) AA = 7.8%
(c) AA = 4.7%
Figure 3: Validation of MoE hyperparameters scaling laws across different activation ratios (AA). “Near-optimal” refers to hyperparameters achieving a loss within 0.25% of the optimal ones.
To validate the generalizability of these laws, we conduct experiments on MoE models with varying activation ratios. We used the derived laws to predict optimal hyperparameters at a compute budget of $3e20$ FLOPs, after fitting them on data up to $1e20$ FLOPs. As shown in Figure 3, the predicted optimal regions effectively capture the best-performing hyperparameters for activation ratios from 4.7% to 10.9%, demonstrating that the laws can be applied to MoE models within this range of activation rates. This confirms that our hyperparameter scaling laws provide a reliable foundation for exploring diverse MoE architectures under fair and near-optimal training conditions.
@@ 2.3 Scaling laws for MoE Optimal Model-Data Allocation
MoE $M^{\text{opt}}=0.1915\cdot C^{0.5095}$ $D^{\text{opt}}=5.2232\cdot C^{0.4905}$
@@ 3 Efficiency Leverage: Metric for Quantifying MoE Compute-Efficiency
$\text{s.t.}\quad\left|\mathcal{L}(C_{\text{moe}};\mathcal{X}_{\text{MoE}})-\mathcal{L}(C_{\text{dense}};\mathcal{X}_{\text{Dense}})\right|$ $\leq\epsilon\quad(\epsilon\to 0)$
@@
To make the analysis tractable, we assume the effects of these dimensions are largely independent and conduct systematic ablation studies. We start with a baseline MoE architecture (2 of 64 experts activated, plus one shared expert) and vary one dimension at a time across a range of compute budgets (from $3e18$ to $3e20$ FLOPs). To ensure a fair and robust comparison, we leverage the findings from our preliminary studies (Sections 2.2 and 2.3). For each architecture and compute budget, we determine the reasonable model size ($M$) and data size ($D$) using our derived allocation laws and configure training with optimal hyperparameters from our hyperparameter scaling laws. This rigorous protocol ensures that each architecture is evaluated at or near its peak potential for a given budget, yielding reliable and cost-effective conclusions. Further details on the experimental setup are provided in Appendix B. Next, we first empirically analyze the impact of each dimension on EL, and then integrate our empirical findings into a unified scaling law to models the relationship between MoE configurations and the resulting EL.
@@ 4.1.1 Optimal Expert Activation Ratio
The IsoFLOPs curves, presented in Figure 5(a), reveal a clear and consistent trend. Across all tested FLOPs budgets (from $1e18$ to $3e20$), loss monotonically decreases with activation ratio, following a power-law pattern. For all configurations, the lowest tested ratio of 0.8% consistently yields the minimum loss. This finding suggests a core principle: for a fixed computational cost, greater model sparsity (i.e., lower activation ratio) leads to higher parameter efficiency.
@@
(a) IsoFLOPs curves for varying AA
(b) Loss and EL scaling curve over varying AA.
Figure 5: Impact of the Activation Ratio AA on Loss and Efficiency. (a) At any fixed compute budget (each colored line), lower activation ratios yield lower loss. The orange stars mark the optimal (lowest) loss point. (b) Loss and EL scaling curves illustrate that EL increases with both higher compute budgets and lower activation ratios, showing that MoE advantages are magnified at scale.
@@ 4.1.3 Optimal Shared Expert Ratio
where $S^{\text{opt}}$ is the optimal shared expert that minimizes the training loss $\mathcal{L}$ under a fixed FLOPs budget $C$, model size $M$, activation ratio $A$, and granularity $G$. Our experiments, as depicted in Figure 7(a), reveal a U-shaped relationship between the shared expert ratio and training loss. The minimum loss is generally achieved at a relatively low shared expert ratio, while having no shared experts ($S=0\%$) usually results in suboptimal performance. Furthermore, we observe a subtle trend where the optimal sharing ratio appears to scale with the compute budget. This is supported by our empirical scaling law (EL) analysis in Figure 7(b), which shows that lower FLOPs budgets ($\leq 10^{20}$) benefit from a slightly higher sharing ratio ($S=16.7\%$), whereas larger budgets ($10^{20}$) achieve greater efficiency with a lower ratio ($S=8.3\%$).
@@
Figure 7: Impact of the Shared Ratio $S$ on Loss and Efficiency. (a) Loss curves demonstrate that a low, non-zero sharing ratio minimizes training loss, outperforming both no shared experts ($S=0\%$) and highly shared configurations.. (b) EL analysis reveal that the optimal sharing ratio is higher ($S=16.7\%$) for smaller FLOPs ($10^{20}$) and decreases to $S=8.3\%$ for larger FLOPs ($10^{20}$).
@@ 4.1.4 Other Configurations of MoE Architecture
First, we analyze replacing the initial MoE layers with dense layers while keeping total FLOPs constant (e.g., 60-layer models with the first 1–3 layers set dense). The experimental results show that replacing the first few layers with dense layers has a minor impact on performance, with efficiency leverage close to 1 within a FLOPs budget of up to 3e203e{20} FLOPs. This adjustment reduces the total number of parameters and mitigates routing imbalances, making it a valuable design optimization. As FLOPs budgets increase, the optimal dense proportion also grows; for example, at $1e18$ FLOPs, the optimal dense proportion is zero. As the compute budget increases to $3e20$ FLOPs, the optimal dense layer proportion shifts to approximately $2/60$ or $3/60$.
@@ 4.2.1 Separable Scaling Laws for Efficiency Leverage
(a) Scaling with Activation Ratio (AA)
@@
Figure 8: Scaling Behavior of Efficiency Leverage (EL). (a) With fixed granularity ($G=2$), EL follows a power law with respect to AA across all tested compute budgets ($C$). (b) With a fixed activation ratio ($A=3.1\%$), EL’s scaling with $G$ conforms to a log-polynomial law across all compute budgets. (c) With both activation ratio (AA) and granularity ($G$) held constant, EL scales with compute according to a standard power law.
@@ 5 Ling-mini-beta: More Efficient MoE Language Model
Based on the findings in Section 4, we identify the efficient architectural configuration within the current MoE framework. To validate the effectiveness of this configuration, we design an MoE model with 0.855 billion active parameters out of a total of 17.5 billion (referred to as “Ling-mini-beta,” a pilot model for Ling-2.0 series) and test it using 1 trillion tokens of training data. Ling-mini-beta is configured with a granularity $G$ of 12 and an activation ratio $A$ of just 3.4%. Referring to Figure 1, at the $1e22$ FLOPs compute budget, we hypothesize that Ling-mini-beta achieves more than 7× in compute-efficiency leverage over a comparable dense model. Concurrently, we train a traditional dense model with 6.1 billion parameters (named “Dense-6.1B”) for comparison. This section presents a detailed analysis of the performance differences between Ling-mini-beta and the conventional dense model Dense-6.1B, highlighting that the active parameter count, training costs, and downstream inference costs of Dense-6.1B are more than seven times those of Ling-mini-beta.
@@ 5.1 Model and Training Details
  • Training Setting. Both Ling-mini-beta and Dense-6.1B were trained using the AdamW optimizer (Loshchilov and Hutter, 2017) with hyperparameters set as follows: $\beta_{1}=0.9$, $\beta_{2}=0.95$, and weight decay of 0.1. Gradient clipping norm is set to 1.0. The learning rate schedule employs a WSD (warmup-stable-decay) strategy (Hu et al., 2024). According to the hyperparameter scaling laws for dense and MoE models, the maximum learning rates were set to 3.78e43.78e{-4} for Ling-mini-beta and 2.93e42.93e{-4} for Dense-6.1B. The batch sizes were configured as 1792 and 2048, respectively.
@@
Model $n_{layers}$ $d_{model}$ $d_{ffn}$ $d_{expert}$ $n_{heads}$ $n_{kv\_head}$ $E$ $E_{a}$ $E_{s}$ NN $N_{a}$
@@ Appendix A Notation
NN Number of non-vocabulary parameters.
@@ Appendix F List of Experimental Models
nlayersn{layers} $d_{model}$ $n_{\text{heads}}$ $E$ $E_{a}$ $E_{s}$ $d_{\text{expert}}$ $B$ $\eta$ Max training FLOPs